Status Report

NOAA GOES SATOPS Morning Report 14 Mar 2001

By SpaceRef Editor
March 14, 2001
Filed under ,

SATOPS MORNING REPORT – March 14, 2001

 

POES

No change in the status of the POES spacecraft.

NOAA-14 rev 31963 / W at 1403Z on March 13:  GAC2B received noisy
from spacecraft with 99.2% of the expected data recovered.  Data not
recoverable due to recorder management.

NOAA-14 rev 31968 / W at 2211Z on March 13:  GAC1B received noisy
from spacecraft with 99.4% of the expected data recovered.  Additional
recovery attempts not requested/attempted.

NOAA-11 rev 64295 / W  at 0154Z on March 14:  The SBUV was
observed to be operating within the limited range of -451 to -590. 
Post pass investigation revealed that it began doing so at 2247Z on March
13.  Dong Han, Jeff Devine, and Tina Baucom notified.  No further
action required as SBUV returned to normal range at 0513Z and appears nominal.

POES operations were nominal over the past 24 hours.

 

DMSP

No change in the status of the DMSP spacecraft.

DMSP operations were nominal over the past 24 hours.

 

GOES

No change in the status of the GOES spacecraft.

All eclipse operations nominal.

At 1225Z on March 13, a GOES-10 frame break was observed and attributed
to high BER (unknown)

At 1939Z on March 13, a GOES-10 frame break was observed and attributed
to thunderstorm activity at CDA.

To review the eclipse schedules for the GOES-8 and GOES-10 spacecraft,
refer to: http://www.ssd.noaa.gov/SSD/GOES/eclipse.html to identify images
canceled to satisfy eclipse or instrument keep-out-zone (KOZ) requirements.

GOES operations were nominal over the past 24 hours.

 

Questions or comments to Tim.Walsh@noaa.gov

SpaceRef staff editor.